# Safe Multi-Agent Reinforcement Learning for Behavior-Based Cooperative Navigation

### v2

@@ V-F Generalizing to More Robots
Finally, we show that by relying only on information from two neighbors, our method can scale to more robots while maintaining a fixed observation size, regardless of the number of robots.
We evaluated the performance of the trained policy with up to eight robots, as shown in Fig. 8. For each robot, we defined two neighbors and tested the policy using the previously introduced goal reaching scenarios. Metrics included the number of goals reached, collisions, and timeouts, with results summarized in Table IV. The robots successfully cooperated to move their centroid to the target locations, with zero collisions. However, as the number of robots increased, we observed an increase in timeouts, attributed to the conservative nature of the MPC filter. These trade-offs are expected in safety-critical systems, where eliminating collisions is paramount.

### v3

@@ V-F Generalizing to More Robots
Finally, we show that by relying only on information from two neighbors, our method scales to more robots while maintaining a fixed observation size, regardless of the number of robots.
We evaluated the performance of the trained policy with up to eight robots, as shown in Fig. 8. For each robot, we defined two neighbors and tested the policy using the previously introduced goal-reaching scenarios. Metrics included the number of goals reached, collisions, and timeouts, with results summarized in Table V. The robots successfully cooperated to move their centroid to the target locations, with zero collisions. However, as the number of robots increased, we observed an increase in timeouts, attributed to the conservative nature of the MPC filter and out-of-distribution scenarios introduced by additional robots. The policy was trained in an environment with only three robots and static obstacles, without exposure to dynamic obstacles. With more robots, they act as dynamic obstacles, leading to local congestion where robots block each other’s paths, ultimately resulting in timeouts. These trade-offs are expected in safety-critical systems, where eliminating collisions is paramount. Addressing these limitations by incorporating dynamic obstacle scenarios into the training, along with integrating an oracle model to predict the future locations of nearby moving agents, represents a promising direction for future work to enhance the system’s robustness and scalability.